# Robust Implicit Adaptive Low Rank Time-Stepping Methods for Matrix Differential Equations

### v3

@@ 4.1 Solid body rotation
$\rho(0,x_{1},x_{2})=e^{-\left(\frac{x_{1}}{0.3}\right)^{2}}e^{-\left(\frac{x_{2}}{0.1}\right)^{2}}$
@@
Figure 2: Displayed is the rank for the Merge, Merge-adapt and Implicit Euler method for the problem with solid body rotation problem. Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@ 4.2 Rotation with anisotropic diffusion
$\rho(0,x_{1},x_{2})=e^{-\left(\frac{x_{1}}{0.3}\right)^{2}}e^{-\left(\frac{x_{2}}{0.1}\right)^{2}}$
@@
In Table 2 we display the errors and rates of convergence for the Merge, Merge-adapt and the implicit Euler method. All three methods behave almost identically (the errors and rates of convergence only differ in the third or fourth digit). However, the time it takes to compute the solution is quite different. In Table 3 we display the CPU times for the different methods. The computations were performed in Matlab on a MacBook Pro with an M2 chip and 16GB RAM. Reasonable effort went into optimizing the codes but the timings should be seen as indicative rather than decisive. As can be seen the growth in CPU time as a function of number of degrees of freedom is very mild for the Merge and Merge-adapt methods in comparison to the classic implicit Euler method. We further note that the Merge-adapt method has comparable cost with the Merge method in this example because the equation involves relatively complex dynamics, which means the explicit prediction is quite costly because of the large number $s.$
$m_{1}=m_{2}$
M time$[s]$
MA time$[s]$
IE time$[s]$
99
0.9
0.9
5.0
199
1.4
1.5
27.1
799
12.8
11.4
926.2
Table 3: Displayed are the CPU times in seconds for the rotation with anisotropic diffusion example. The results are all for $n_{T}=320$. As can be seen the growth in CPU time as a function of number of degrees of freedom is very mild for the Merge and Merge-adapt methods in comparison to the classic implicit Euler method.
@@
Figure 3: Displayed is the rank for the Merge, Merge-adapt and Implicit Euler method for the problem with rotation and anisotropic diffusion. Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@ 4.3 Anisotropic diffusion
Figure 4: Displayed is the rank for the Merge, Merge-adapt and Implicit Euler method for the problem with anisotropic diffusion initial data $\sin(\pi x_{1})\sin(\pi x_{2}),$ (left) and $\sin(2\pi x_{1})\sin(2\pi x_{2})$ (right) and the solid body rotation problem (right). Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@
4.4 Solid body rotation with isotropic diffusion
$m_{1}=m_{2}$
time MA [s]
time M [s]
$\Delta t/h$ : $\mu\Delta t/h^{2}$
999
1.9 [0.1%]
3.1
0.39 : 0.020
1999
3.5 [0.1%]
6.0
0.79 : 0.79
3999
7.1 [0.4%]
15.5
1.57 : 0.31
7999
52.8 [72%]
46.2
3.14 : 1.26
Table 6: CPU-times for the problem with solid body rotation with isotropic diffusion. The numbers in brackets are the percentages for how often MA has a residual that necessitates the addition of the BUG space.
In this experiment we time a problem with solid body rotation and isotropic diffusion, corresponding to, $r_{1}(x_{1})=x_{1}$ $r_{2}(x_{2})=-x_{2}$, $a_{1}=b_{1}=a_{4}=b_{4}=10^{-4}$, and $a_{j}=b_{j}=0,\,j=2,\ldots,3$. We focus on the comparison of the computational cost of Merge and Merge-adapt schemes for large mesh size. We start from the rank-1 initial data
$\rho(0,x_{1},x_{2})=e^{-\left(\frac{x_{1}}{0.3}\right)^{2}}e^{-\left(\frac{x_{2}}{0.1}\right)^{2}}$
and evolve the solution until time $t=0.25\pi$ using a base timestep of $\Delta t=0.25\pi/n_{T}$, with $n_{T}=1000$. We carry out the computation on very fine grids with $m_{1}=m_{2}=999,1999,3999,7999$. Here the size of the diffusion is chosen so that the ratios $\Delta t/h$ : $\mu\Delta t/h^{2}$ change from smaller than one to larger than one. We report the CPU times for the M and MA methods in Table 6. As can be seen the MA is faster than the M method when the number of timesteps where the BUG space is used is small. The more intensive use of the BUG space appears to happen when the ratio $\mu\Delta t/h^{2}$ increases above one.

### v4

@@ 4.1 Solid body rotation
$\rho(0,x_{1},x_{2})=e^{-\left(\frac{x_{1}}{0.3}\right)^{2}}e^{-\left(\frac{x_{2}}{0.1}\right)^{2}},$
@@
Figure 2: Displayed is the truncated rank for the Merge, Merge-adapt and Implicit Euler method for the problem with solid body rotation problem. Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@ 4.2 Rotation with anisotropic diffusion
$\rho(0,x_{1},x_{2})=e^{-\left(\frac{x_{1}}{0.3}\right)^{2}}e^{-\left(\frac{x_{2}}{0.1}\right)^{2}},$
@@
In Table 2 we display the errors and rates of convergence for the Merge, Merge-adapt and the implicit Euler method. All three methods behave almost identically (the errors and rates of convergence only differ in the third or fourth digit). However, the time it takes to compute the solution is quite different.
@@
Figure 3: Displayed is the truncated rank for the Merge, Merge-adapt and Implicit Euler method for the problem with rotation and anisotropic diffusion. Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@ 4.3 Anisotropic diffusion
Figure 4: Displayed is the truncated rank for the Merge, Merge-adapt and Implicit Euler method for the problem with anisotropic diffusion initial data $\sin(\pi x_{1})\sin(\pi x_{2}),$ (left) and $\sin(2\pi x_{1})\sin(2\pi x_{2})$ (right) and the solid body rotation problem (right). Here “Fail for MA” is the indicator where the BUG space is needed for the Merge-adapt method. For the Implicit Euler method we constantly use the threshold $\Delta t^{2}$ when computing the rank via the truncated SVD. All the computations are done with $m_{1}=m_{2}=799$ and $n_{T}=320.$
@@
4.4 Timing results
$m_{1}=m_{2}$ time MA time M time MA time M time MA time M time IE
499 2 [0.1%] 18 6 [0.0%] 19 10 [45%] 28 1315
999 4 [0.1%] 68 12 [0.0%] 61 38 [48%] 260 153800
1999 11 [0.1%] 244 37 [0.0%] 129 535 [51%] 1645 N/A
Table 5: CPU-times (in seconds) for the problems 4.1 4.2 and 4.3 from lef to right. The numbers in brackets are the percentages for how often MA has a residual that necessitates the addition of the BUG space.
In this experiment we repeat the experiments in Sections 4.1 4.2 and 4.3. We use $n_{T}=1000$ and carry out computations on fine grids with $m_{1}=m_{2}=499,999,1999$. We report the CPU times for the M and MA methods in Table 5. As can be seen the MA is generally faster than the M method. The more intensive use of the BUG space appears to happen when the ratio $\mu\Delta t/h^{2}$ is large. For the example in Section 4.3 we have also added the time needed for the IE method for the two coarsest resolutions (the computation on the finest mesh could not fit on the computer used). For the IE timing we have only measured the time for the forward and backward substitutions in each timestep, not for the initial LU-factorization.
@@ 6 Acknowledgement
Declarations
This research is partially supported by DOE Office of Advanced Scientific Computing Research under the Advanced Research in Quantum Computing program, subcontracted from award 2019-LLNL-SCW-1683, NSF DMS-2208164, NSF DMS-2011838, DOE grant DE-SC0023164 and Virginia Tech. The authors have no relevant financial or non-financial interests to disclose.
Data availability statement
Data sets generated during the current study are available from the corresponding author on reasonable request.